Lina is 5 years older than Jane. Eight years ago she was three times as old as Jane. What is the age of each girl? Please Help...
Answer by scott8148(6628)

"Lina is 5 years older than Jane" ___ L = J + 5

"Eight years ago she was three times as old as Jane" ___ L - 8 = 3(J - 8)